# Multilingual Performance of a Multimodal Artificial Intelligence System on Multisubject Physics Concept Inventories

### v1

@@ IV.1 Scores by inventory and language
Also indicated in Table 7 is the relative performance compared to published scores. For most inventories in most nominal languages, the AI system outperforms published post-instruction student scores at the undergraduate level. It is notable that for each inventory, this performance differential is fairly consistent across nominal languages.
IV.2 Score comparison by subject area
Table 8 shows a comparison of average post-test student scores found in the literature for undergraduate-level courses (see column “%Perc” in Tables 3-6) and average AI scores. We calculate the AI scores based on the scores listed in Table 7, taking the average across all languages. With the exception of laboratory skills (LAB), the AI outperforms average undergraduate students in every subject category. Note that the average AI score would usually be even higher, if we only considered tests in English.
| Category | Avg. Student Score | Avg. AI Score | AI/Students |
| AST | 48 | 79 | 1.64 |
| EM-C | 46 | 58 | 1.25 |
| EM-F | 54 | 63 | 1.17 |
| LAB | 39 | 37 | 0.94 |
| MATH | 60 | 71 | 1.19 |
| MECH | 54 | 67 | 1.24 |
| OPT | 38 | 60 | 1.57 |
| QUANT | 50 | 65 | 1.30 |
| REAS | 58 | 75 | 1.29 |
| RELA | 71 | 75 | 1.06 |
| THERM | 60 | 77 | 1.28 |
IV.3 Performance by subject area and image usage
Table 9 shows the performance by subject area and usage of images, which allows insight into the multimodal capabilities of the AI system. Here, we calculate the performance based on all questions that are in inventories that address the subject area and, for the three columns on the right, fulfill the criteria regarding the presence and use of images. For each question, we coded if it contained or referred to an image, i.e. a visual representation (sketch, graph, diagram, etc.) and if interpreting the image was required for correctly solving the task (required image), or if the image was redundant and the task could be solved by only reading the text of the problem (unneeded image).
Overall, the AI performs best in Astronomy (AST), followed by thermodynamics (THERM), relativity (RELA), and reasoning (REAS). The inventories on Astronomy tend to focus on knowledge, while THERM, RELA, and REAS inventories tend to use exact language. The AI system is particularly weak in laboratory skills (LAB), which include strategies for data collection and analysis, Optics (OPT), mechanics (MECH), and electricity and magnetism (both circuits (EM-C) and fields (EM-F)).

### v3

@@ IV.3 How does the AI’s performance compare to student performance at the undergraduate level?
In Table 6 we use color to indicate the relative performance of GPT-4o compared to published post-instruction scores at the undergraduate student level found in the literature (see column “%Post” in Tables 2-5). For most inventories in most languages, the AI system outperforms the student average. Where student data was available, the AI outperformed average undergraduate post-instruction scores in 68.9% of cases. To obtain this value, we compared student averages on an inventory to AI’s performance on the inventory for each available language and summed across all inventories.
As we have pointed out in the Introduction (Section I.4), the comparison of average performance of GPT-4o to that of students provides a rough proxy measure for its capabilities in solving physics conceptual tasks in relation to student capabilities. However, caution is needed when interpreting these results, as the profile of GPT-4o’s and the ”average” student’s strengths and difficulties can differ.
Figure 5 shows the distributions of post-test student scores and AI scores across all languages, grouped by subject category. The averages are indicated by vertical markers. With the exception of Laboratory skills (LAB), GPT-4o outperforms the average undergraduate student in every subject category of concept inventories, with the most significant differences in Astronomy and Reasoning. The extremely wide distribution of the mechanics (MECH) scores is mostly due to the fact that the Force Concept Inventory (FCI) is available in a wide variety of languages, including languages such as Punjabi and Tamil, which GPT-4o does not appear to adequately master (see Fig. 4). The wide distributions in Quantum Physics and Optics are again due to inventories with mostly graphical, visual representations.
Well-designed assessment instruments for teaching typically include multiple-choice distractors that probe specific student misconceptions. Since we drew our questions from the PhysPort library, we can assume that most concept inventories in our dataset were designed carefully. Therefore, we expect students to gravitate toward certain incorrect answers that reflect common misconceptions. This raises the question of whether the AI similarly gravitates toward specific incorrect answers. Although not all the assessments analyzed in this study had five possible choices, the majority did, allowing us to use this as a reference point (see Appendix C and Table 10 for more detail). If the AI were selecting incorrect answers randomly, we would expect a 75% probability that two incorrect answers differ and a 25% chance that they would be the same. However, for our data, this was reversed: 66% of all items had the same two incorrect answers, while 34% had different ones. The same pattern held for cases where all three answers were incorrect. In a random scenario,we would expect 6.25% of items to have three identical answers, 37.5% to have three different ones, and 56.25% to have two the same (see Appendix C for more details). In reality, AI responses were far from random and gravitate toward specific incorrect answers: 53% of such items had three identical incorrect answers, while only 8% had three different ones. Similar trends were observed for both English and non-English items. This analysis suggests that the AI system exhibits some consistency in its choice of incorrect answers. However, further research is necessary to investigate how students and AI systems may differ in how they respond to different types of multiple-choice options.
IV.4 How does the presence of images influence the performance of the AI system?
Here, we calculate the performance based on the sum across individual inventory items appearing in the different inventories (all languages combined). For each item, we coded if it contained or referred to an image — that is, a visual representation such as a sketch, graph, diagram — and whether interpreting the image was required for correctly solving the task (“required image”), or if the image was redundant, meaning all information required for solving the task was already provided in the text (“unneeded image”). The percentages shown in Figure 6 were obtained by dividing the number of correct responses in each category with the total number of submissions in that category.
Overall, we found that the performance on text-only tasks was 81% , compared to 79% on tasks containing unneeded images, and just 49% on tasks with required images (see Figure 6).